STUDY OF HIGHLY-PREDISSOCIATIVE STATES OF FREE RADICALS USING TWO-COLOR RESO-NANT FOUR-WAVE MIXING TECHNIQUE

\underline{YUAN-PERN LEE}, Department of Chemistry, National Tsing Hua University, Hsinchu, Taiwan 30043.

We demonstrate an application of two-color resonant four-wave mixing (TC-RFWM) spectroscopy to detect highly predissociative levels of the C ^{2}Σ^{+} AND D ^{2}Π states of CH in a hostile environment of an oxyacetylene flame. The 1-1 band of the C - X transition is detected with the probe and the grating wavelengths in resonance with the A ^{2}Δ - X ^{2}Π and C ^{2}Σ^{+} - X ^{2}Π transitions, respectively. Six branches of the C - X (1-1) system are spectrally resolved for the first time; in total 124 lines detected in this work correspond to excitation of the C ^{2}Σ^{+} (v = 1) state up to N′ = 23. Observed wave numbers are fitted to yield improved spectral parameters of the C ^{2}Σ^{+} state. The D state was observed by using two grating (probe) beams in resonance with the D ^{2}Π - B ^{2}Σ^{−} transition and the pump beam in resonance with a selected B ^{2}Σ^{−} - X ^{2}Π transition. A total of 86 lines associated with transitions to the D ^{2}Π (v = 0) state with rotational quantum number N′ up to 16 were detected; parameters for centrifugal-distortion (D, H), spin-rotation interaction (γ), and Λ-doubling (p, q, q_{0}) of the D ^{2}Π (v = 0) state were deduced for the first time. The predissociative line width is about 2 cm^{−1} for N′ = 1 and gradually increases to 7 cm^{−1} for N′ = 15. Predissociation mechanisms of both the C and the D states are discussed. If time permits, other systems will be briefly described.